OpenStudy (anonymous):

What value of x makes the equation 3(x-6)-8x= -2+5(2x+1) true? Show your work please.

OpenStudy (mustbethestars):

Okay do you know what to do first?

OpenStudy (mustbethestars):

Come on man I can't give you a direct answer work with me here bud.

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

sorry! my internet stopped!

OpenStudy (mustbethestars):

Oh my bad. You have to distribute..... so distribute 3(x-6) and 5(2x+1)

OpenStudy (anonymous):

okay

OpenStudy (mustbethestars):

Okay let's see your work :)

OpenStudy (anonymous):

alright im going to do it on paper first

OpenStudy (mustbethestars):

Okay :3

OpenStudy (anonymous):

how would i distribute 3 into x?

OpenStudy (anonymous):

3x?

OpenStudy (mustbethestars):

Ye. 3x and then 3*6 which is 18 so 3x-18...... now do the next one

OpenStudy (anonymous):

okay

OpenStudy (mustbethestars):

Let's see your work :3

OpenStudy (anonymous):

would it be 7x (+1)?

OpenStudy (anonymous):

if so it'd be 8x. right?

OpenStudy (mustbethestars):

No what is 5 * 2x?

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

10x?

OpenStudy (mustbethestars):

Correct and 5 *1 is 5 so it is 10x+5..... So now we got: 3x-18-8x = 10x + 5 - 2

OpenStudy (mustbethestars):

And we know 5-2 is 3. 3x-18-8x = 10x + 3 Now what do we do?

OpenStudy (anonymous):

im not sure....prove what makes it true...?

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

would that be 3?

OpenStudy (mustbethestars):

No. 3x-18-8x(+8x) = 10x (+8x) + 3 3x - 18 = 18x + 3

OpenStudy (mustbethestars):

Wait I think I messed up .-.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i think the first part of it looks weird

OpenStudy (raffle_snaffle):

Lets solve for x and redo this problem

OpenStudy (anonymous):

okay

OpenStudy (raffle_snaffle):

3(x-6)-8x= -2+5(2x+1) lets start with the left side what is 3*(x-6) ?

OpenStudy (raffle_snaffle):

distribute the 3 through